# Inverse-Squeezing Kennedy Receiver for Near-Helstrom Discrimination of Displaced-Squeezed BPSK

### v1

@@ Abstract
To address the discrimination problem of binary phase-shift keyed displaced squeezed vacuum states (S-BPSK), this paper proposes an Inverse-squeezing Kennedy (IS-Kennedy) receiver. This architecture incorporates an inverse-squeezing operator following the displacement operation of a conventional Kennedy receiver, mapping the S-BPSK signals onto equivalent large-amplitude coherent states. Furthermore, it employs a photon-number-resolving (PNR) detector to perform maximum a posteriori (MAP) decision-making. Theoretical analysis demonstrates that, under ideal conditions, the IS-Kennedy receiver effectively translates the transmitter’s squeezing resources into a displacement gain at the receiver. Consequently, its error probability approaches the Helstrom bound across the entire energy spectrum, remaining within a constant factor of 3 dB. In the low-photon-number regime ($N\approx 0.6$), the proposed scheme surpasses the coherent-state limit, achieving an error rate below 1%. Furthermore, this paper provides an in-depth analysis of system performance under non-ideal conditions, revealing the robustness of PNR detection against background dark counts and a characteristic “parity photon-number step” saturation effect arising from squeezing parameter mismatch.
@@ I Introduction
As an important topic in the field of quantum information, quantum-states discrimination directly affects the performance of quantum communication systems. To date, we have consistently regarded coherent states (CSs) as carriers for quantum communication and have proposed numerous excellent discrimination schemes, such as the Kennedy receiver[15], the optimized displacement receiver[22, 26], multi-stage receiver[23, 14, 4, 18], the conditional pulse nulling receiver[27], and hybrid receiver [17, 1, 19, 28]. This is primarily because coherent states retain their coherence after passing through lossy bosonic channels, demonstrating that they can maintain state purity in most optical transmission media, including but not limited to free space and optical fibers.[9, 25, 16]. However, coherent states have inherent discrimination limitations: any two coherent states are non-orthogonal and exhibit overlap in phase space. Since the Helstrom Bound (HB) is closely related to the degree of overlap between quantum states[10, 11], coherent states have a theoretical lower bound on discrimination error probability.
In recent years, quantum-state discrimination schemes based on non-Poissonian photon statistics have attracted widespread attention[20, 13, 21, 5, 3, 24]. For example, Giovanni Chesi and Stefano Olivares proposed using squeezed vacuum states ($S(r)\left|0\right\rangle$ where $S(r),r\in\mathbf{R}_{+}$ is squeeze operator and $\left|0\right\rangle$ is vacuum state) as information carriers to construct a binary communication channel[5]. By applying a displacement operation $D(\pm\alpha)$ to the squeezed vacuum state, two phase-opposite displaced squeezed vacuum states (DSSs) $\left|\pm\alpha,r\right\rangle$ can be generated. Then they encoded BPSK symbols into two phase-opposite DSSs (S-BPSK):
$\left\{\begin{aligned} \text{symbol}\ ^{\prime}0^{\prime}:\ &\left|\psi_{0}\right\rangle=D(-\alpha)S(r)\left|0\right\rangle=\left|-\alpha,r\right\rangle\\ \text{symbol}\ ^{\prime}1^{\prime}:\ &\left|\psi_{1}\right\rangle=D(+\alpha)S(r)\left|0\right\rangle=\left|+\alpha,r\right\rangle\end{aligned}\right$. (1)
This scheme significantly reduces the overlap between the carrier states by squeezing its position quadrature $\hat{X}$. Theoretical studies show that when the communication channel energy $N$ is fixed, not only the HB for S-BPSK can be significantly lower than that for BPSK coherent states (C-BPSK), but the error probability of homodyne detection (Standard Quantum Limit, SQL) for S-BPSK can even surpass the HB for C-BPSK at a low mean photon number of $N\approx 0.659$ [5].
Unlike the work of Giovanni Chesi and Stefano Olivares, this paper proposes a quantum-enhanced receiver for discriminating the S-BPSK signal, which is called as the Inverse-squeezing Kennedy Receiver (IS-Kennedy). The IS-Kennedy combines Gaussian operations (displacement and squeeze) and photon-number-resolving detection to approach the Helstrom bound for S-BPSK within a constant factor, while outperforming coherent-state receivers in the low-energy regime. Theoretical analysis shows that, for the same input mean photon number, the IS-Kennedy can achieve an error probability lower than the Helstrom bound for C-BPSK channels in the low-photon-number regime. For example, when $N\approx 0.6$ the error probability falls below 1% in our simulations.
The structure of this paper is as follows: Section II reviews the fundamental theory of S-BPSK discrimination; Section III provides a detailed analysis of the principle and performance of the IS-Kennedy; Section IV explores the performance of IS-Kennedy with an imperfect PNR detector (quantum efficiency $\eta$ and dark count rate $\nu$); Section V explores the performance of IS-Kennedy under inverse-squeezing parameter mismatch; finally, the paper is summarized, and potential application scenarios are discussed.
II Channel model and performance benchmarks
In this section we summarize the signal model and the fundamental performance limits for binary phase-shift keyed (BPSK) communication based on displaced squeezed vacuum states, and we fix the notation used in the rest of the paper.
For binary discrimination between quantum states $\left|\psi_{0}\right\rangle$ and $\left|\psi_{1}\right\rangle$ with prior probabilities $\pi_{0}=\pi_{1}=\frac{1}{2}$, the minimum error probability (Helstrom bound) is given by [10, 5]:
$P_{\text{HB}}^{\text{DSS}}\left(\alpha,r\right)$ $=\frac{1}{2}\left(1-\sqrt{1-\left|\left\langle\psi_{0}|\psi_{1}\right\rangle\right|^{2}}\right)$ (2)
$=\frac{1}{2}\left(1-\sqrt{1-\exp\left(-4\alpha^{2}\text{e}^{2r}\right)}\right)$
As a practically relevant benchmark we also consider homodyne detection along the X quadrature, which attains the standard quantum limit (SQL) for BPSK under ideal conditions. Given the input is $\rho_{i}=\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|$, the conditional probability density of the homodyne outcome $x$ ,
$P\left(x|\rho_{i}\right)=\int_{-\infty}^{+\infty}\text{d}p\cdot W\left(x,p|\rho_{i}\right)$ (3)
$=\frac{\text{e}^{r}}{\sqrt{\pi}}\,\exp\left\{-\,\text{e}^{2r}\left[x-(-1)^{i+1}\sqrt{2}\alpha\right]^{2}\right\},\ i\in{0,1},$
which $W\left(x,p|\rho_{i}\right)$ is the Wigner function of $\rho_{i}$ (see Appendix). And the resulting error probability is
$P_{\text{SQL}}^{\text{DSS}}=\int_{0}^{+\infty}P\left(x|\rho_{0}\right)\text{d}x=\frac{1}{2}\text{erfc}\left(\sqrt{2}\alpha\text{e}^{r}\right)$ (4)
To parameterize the contribution of squeezing, we introduce the squeezing fraction
$\beta\triangleq\frac{\sinh^{2}(r)}{N}$ (5)
which quantifies the share of the total energy ($N=\left|\alpha\right|^{2}+\sinh^{2}(r)$) stored in squeezing [5]. Therefore, the parameter $\alpha$ and $r$ can be expressed as
$\left\{\begin{aligned} &\alpha=\sqrt{N\left(1-\beta\right)},\quad r=\sinh^{-1}\sqrt{N\beta}\\ &\alpha\text{e}^{r}=\sqrt{N\left(1-\beta\right)}\left(\sqrt{1+N\beta}+\sqrt{N\beta}\right)\end{aligned}\right$. (6)
Accordingly, both the Helstrom bound and the SQL for S-BPSK become functions of $\left(N,\beta\right)$, denoted $P_{\text{SQL}}^{\text{DSS}}\left(N,\beta\right)$, and $P_{\text{HB}}^{\text{DSS}}\left(N,\beta\right)$. The squeezing fraction can be optimized at fixed total energy. By minimizing $P_{\text{HB}}^{\text{DSS}}\left(N,\beta\right)$ over $\beta$ one obtains the optimal value [5]:
$\beta_{\text{opt}}(N)=\frac{N}{2N+1},\quad\alpha\text{e}^{r}=\sqrt{N(N+1)}$ (7)
Figure 1: The ratios of $P_{\text{HB}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)$, $P_{\text{SQL}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)$ and $P_{\text{SQL}}^{\text{CS}}\left(N\right)$ to $P_{\text{HB}}^{\text{CS}}\left(N\right)$ vs. input state energy N. Here, $P_{\text{SQL}}^{\text{CS}}\left(N\right)=\frac{1}{2}\left[1-\text{erf}\left(\sqrt{2\cdot N}\right)\right]$ and $P_{\text{HB}}^{\text{CS}}\left(N\right)=\frac{1}{2}\left[1-\sqrt{1-\exp\left(-4\cdot N\right)}\right]$.
Both the Helstrom bound and the SQL for S-BPSK are rewritten as:
$P_{\text{SQL}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\text{erfc}\left(\sqrt{2N\left(N+1\right)}\right)$ (8)
$P_{\text{HB}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\left(1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}\right)$ (9)
When the DSSs are prepared with this choice, the S-BPSK Helstrom bound is significantly lower than that of C-BPSK at the same energy. In particular, for $N\approx 1$ the S-BPSK Helstrom bound already reaches the order of $10^{-4}$, whereas C-BPSK requires approximately twice the energy to achieve a comparable error probability. Moreover, in the low-photon number regime the S-BPSK SQL can fall below the Helstrom bound of C-BPSK, with the crossover occurring around $N\approx 0.659$ (see Fig. 1). At $N=1.0$, the HB for S-BPSK surpasses the SQL for C-BPSK by a remarkable 17.39dB, and the SQL for S-BPSK also is 2.94dB lower. In the subsequent analyses, all DSSs employ the optimal squeezing fraction $\beta_{\text{opt}}$.
III Inverse-squeezing Kennedy Receiver under Ideal Conditions
In this section, we propose a novel receiver architecture for S-BPSK discrimination: the Inverse-squeezing Kennedy (IS-Kennedy) receiver. The schematic structure is illustrated in Fig. 2. Compared to the conventional Kennedy receiver, the IS-Kennedy incorporates two key modifications: (i) the insertion of an inverse-squeezing (IS) module after the displacement operation, which applies the operator S(r)S(-r) to the signal; and (ii) the upgrade of the terminal single-photon detector (SPD) to a photon-number-resolving detector (PNRD), enabling maximum a posteriori (MAP) decision-making under generalized statistical models. The squeezing amplitude of the IS module is matched to that of the transmitter, while the squeezing axis is rotated by $\pi/2$ (orthogonal) relative to the transmitter. This configuration cancels the initial squeezing and produces an equivalent amplification of the signal’s mean photon number. Effectively, the IS-Kennedy converts the transmitter’s squeezing resources into an equivalent amplitude gain at the receiver, thereby reducing the S-BPSK discrimination problem to an “amplified Kennedy nulling problem.” Furthermore, the PNRD provides the necessary statistical degrees of freedom for optimal MAP decisions under non-ideal conditions, which will be discussed later.
@@ III Inverse-squeezing Kennedy Receiver under Ideal Conditions
Subsequently, $\sigma_{i}$ passes through the IS module, mapping the S-OOK alphabet to a coherent-state OOK (C-OOK) alphabet $\zeta_{i}$:
@@
In the general case, $\gamma=\alpha\cosh r+\alpha^{*}\sinh r$. Under the condition $\alpha\in\mathbb{R}_{+}$ adopted in this work, $\gamma=\alpha e^{r}$. Therefore, we define:
$\gamma\triangleq\alpha e^{r},\qquad N_{\text{eff}}\triangleq|\gamma|^{2}=|\alpha|^{2}e^{2r}.$ (12)
With the optimal energy allocation $\beta_{\text{opt}}$, the effective photon number can be further expressed as $N_{\text{eff}}=N(N+1)>N$. This indicates that the IS module transforms the transmitter’s squeezing resources into a “coherent-state separation gain” directly accessible by the receiver. This results in a larger Euclidean distance between the two states in phase space (Fig. 3(a-c)) and simultaneously reduces the overlap of the Fock-basis populations of $\zeta_{i}$ (Fig. 3(e-f)).
Figure 3: Phase-space representations (a–c) and Fock-basis populations (d–f) of the input states $\rho_{i}$ in the IS-Kennedy for mean photon number $N=1.0$ and squeezing fraction $\beta=\beta_{\text{opt}}$. Panels (a) and (d) show the input states $\rho_{i}$; panels (b) and (e) show the states after the full IS-Kennedy transformation $\sigma_{i}=D(\alpha)\rho_{i}D^{\dagger}(\alpha)$. Panels (c) and (f) show the states after inverse-squeezing (IS) $\zeta_{i}=U\rho_{i}U^{\dagger}$ with $U=S(-r)D(\alpha)$.
Next, PNR detection is performed on $\zeta_{i}$, yielding an output photon number $n$. The positive operator-valued measure (POVM) for an ideal PNR detector is given by:
@@
From Eq. (11), we have $|\zeta_{0}=\ket{0}\bra{0}$ and $|\zeta_{1}=\ket{2\gamma}\bra{2\gamma}$, with mean photon numbers $\mu_{0}=0$ and $\mu_{1}=|2\gamma|^{2}=4N_{\text{eff}}$, respectively. Thus, the probability of detecting $n$ photons given transmitted symbol $i$ is:
$P(n|i)=\text{Tr}\left(\zeta_{i}\Pi_{n}^{\text{ideal}}\right)=\left\{\begin{aligned} &\text{Poiss}\left(n;\mu_{0}\right)=\delta_{0,n},&i=0,\\ &\text{Poiss}\left(n;\mu_{1}\right),&i=1,\end{aligned}\right$. (14)
where $\text{Poiss}\left(n;\mu\right)=\mathrm{e}^{-\mu}\frac{\mu^{n}}{n!}$. The posterior probabilities are given by:
$P(i=0|n)=\frac{\pi_{0}P(n|i=0)}{\sum_{k}\pi_{k}P(n|i=k)},\quad P(i=1|n)=\frac{\pi_{1}P(n|i=1)}{\sum_{k}\pi_{k}P(n|i=k)}.$ (15)

### v2

# Squeezing Mismatch in the Inverse-Squeezing Kennedy Receiver for Binary Displaced Squeezed-State Discrimination

@@ Abstract
We analyze inverse-squeezing mismatch in the IS-Kennedy receiver for binary displaced squeezed vacuum state discrimination. Mismatch is shown to be equivalent to a residual squeezing after nulling, which modifies the output photon-number statistics and makes the optimal maximum-a-posteriori (MAP) rule generally non-single-threshold. We find that the receiver is much more sensitive to phase mismatch than to amplitude mismatch. Under amplitude-only mismatch, the saturation error with finite-resolution photon-number-resolving detection exhibits a parity-step scaling, decreasing only when the detector resolution crosses even-photon thresholds. These results clarify the physical origin of mismatch-induced degradation and identify phase locking as the key requirement for practical implementations.
@@ I Introduction
Quantum receivers can surpass the standard quantum limit (SQL) in binary optical-state discrimination by exploiting the structure of the signal set and appropriate measurement design. Beyond the extensively studied coherent-state case, displaced squeezed vacuum states provide an important nonclassical modulation format, since squeezing can improve discrimination performance when properly matched to the receiver [1]. For binary phase-shift-keyed displaced squeezed states (S-BPSK), the optimal squeezing fraction minimizes both the Helstrom bound and the SQL, making this alphabet a natural platform for studying squeezing-enhanced quantum communication [7, 2, 8].
Recently, an inverse-squeezing Kennedy (IS-Kennedy) receiver was proposed for S-BPSK discrimination [1]. In the ideal case, the receiver first performs Kennedy nulling and then applies an inverse-squeezing operation, thereby mapping the squeezed on-off-keyed output into a coherent-state on-off-keyed alphabet before photon counting. This transformation converts the squeezing resource into an effective coherent-energy gain and allows the receiver to outperform the S-BPSK SQL [2] over a broad signal range.
In practice, however, the inverse-squeezing stage is highly sensitive to implementation imperfections. Deviations in the squeezing magnitude or phase may arise from optical loss, pump fluctuations, and phase-locking errors [3, 4], and even small mismatch can qualitatively change the receiver output statistics. This practical issue is particularly important because the performance advantage of the IS-Kennedy receiver relies on accurate cancellation of the input squeezing. In this letter, we investigate inverse-squeezing mismatch in the IS-Kennedy receiver for S-BPSK discrimination. We show that mismatch is equivalent to a residual squeezing after nulling, which makes the optimal maximum-a-posteriori (MAP) decision generally non-single-threshold. We further show that the receiver is much more sensitive to phase mismatch than to amplitude mismatch, and that under amplitude-only mismatch the saturation error with finite-resolution PNR detection exhibits a parity-step scaling. These results clarify the physical origin of mismatch-induced degradation and identify the phase-locking accuracy required for practical implementations of inverse-squeezing receivers.
II Receiver modeling
Here, let us consider the modeling of the Inverse-squeezing Kennedy receiver [1]. Fig. 1 shows the schematic of the inverse-squeezing Kennedy receiver.
Figure 1: The schematic of the inverse-squeezing Kennedy receiver.
The S-BPSK signals to be discriminated are defined as [2]:
@@ II Receiver modeling
Without loss of generality, $\alpha$ is chosen to be real, and the mean photon numbers of the signals are $N=|\alpha|^{2}+\sinh^{2}(r)$. Throughout this letter, a priori probability $\pi_{i}$ of these displaced squeezed vacuum states are assumed to be equal for 2, i.e. $\pi_{i}=\frac{1}{2}$.
Reference [2] defines the squeezing fraction $\beta$ (the share of the total energy stored in squeezing) to quantify the advantage of S-BPSK:
$\beta\triangleq\frac{\sinh^{2}(r)}{N}.$ (2)
In addition, under ideal condition, the optimal squeezing fraction is $\beta_{\text{opt}}=\frac{N}{2N+1}$, at which point Helstrom bound and SQL of S-BPSK are minimized [2]. In this letter, all S-BPSK employ the optimal squeezing fraction $\beta_{\text{opt}}$:
$\alpha=\sqrt{N(1-\beta_{\text{opt}})},\quad r=\sinh^{-1}\sqrt{N\beta_{\text{opt}}},\quad\alpha\text{e}^{r}=\sqrt{N(N+1)}$ (3)
II.1 Receiver under ideal condition
Under ideal condition, firstly, the received signal interferes with a local oscillator (LO) on a highly transmissive beam splitter ($\tau\to 1$) to perform the Kennedy nulling displacement $D(\alpha)$. Consequently, the S-BPSK alphabet is mapped to a squeezed OOK (S-OOK) alphabet $\sigma_{i}$:
@@ II.1 Receiver under ideal condition
Subsequently, $\sigma_{i}$ passes through the IS module S(r)S(-r), mapping the S-OOK alphabet to a coherent-state OOK (C-OOK) alphabet $\zeta_{i}$:
@@
In the general case, $\gamma=\alpha\cosh r+\alpha^{*}\sinh r$. In this letter $\alpha\in\mathbb{R}_{+}$, so $\gamma=\alpha e^{r}$. Therefore, we define:
$\gamma\triangleq\alpha e^{r},\quad N_{\text{eff}}\triangleq|\gamma|^{2}=|\alpha|^{2}e^{2r}.$ (6)
Next, PNR detection is performed on $\zeta_{i}$, yielding an output photon number $n$. The positive operator-valued measure (POVM) for an ideal PNR detector with resolution M, denoted as PNR(M), is given by:
$\Pi^{(\text{M})}_{n}=\left\{\begin{aligned} &\Pi_{n},\quad n=0,1,\cdots,\text{M}-1\\ &\mathbb{I}-\sum_{k=0}^{\text{M}-1}\Pi_{k},\quad n=\text{M}\end{aligned}\right$. (7)
@@
Therefore, the probability of detecting $n$ photons given hypothes $^{\prime}i^{\prime}$ is:
$P(n|i)=\text{Tr}\left\{\zeta\cdot\Pi_{n}^{(\text{M})}\right\}=\left\{\begin{aligned} &\text{Poiss}(n;\mu_{i}),\quad n=0,1,\cdots,\text{M}-1\\ &1-\sum_{k=0}^{\text{M}-1}\text{Poiss}(k;\mu_{i}),\quad n=\text{M}\end{aligned}\right$. (8)
where $\text{Poiss}(n;\mu)=\text{e}^{-mu}\frac{\mu^{n}}{n!}$; $\mu_{0}=0$ and $\mu_{1}=4N_{\text{eff}}$ are the mean photon number of $\zeta_{0}$ and $\zeta_{1}$, respectively. Then, the posterior probabilities are given by:
$P(i|n)=\frac{\pi_{i}P(n|i)}{\sum_{k}\pi_{k}P(n|k)}=\frac{P(n|i)}{\sum_{k}P(n|k)}$ (9)